OpenStudy (anonymous):

Need help. Use logarithmic differentiation to find the derivative of the following equation. y = (6x + 1)^5(x4 − 1)^6

OpenStudy (anonymous):

wait it's actually.. y= ((6x+1)^5) * ((x^4)-1)^6

OpenStudy (anonymous):

take the log, simplify using properties of the log, take the derivative of the simplified thing, then multiply by the original function

OpenStudy (turingtest):

\[y = (6x + 1)^5(x^4 − 1)^6\]\[\ln y=5\ln(6x+1)+6\ln(x^4-1)\]can you differentiate this?

OpenStudy (anonymous):

i think you get \[5\ln(6x+1)+6\ln(x^4-1)\] \[5\ln(6x+1)+6(\ln(x+1)+\ln(x-1)+\ln(x^2+1))\] as a first step

OpenStudy (anonymous):

but that may be excessive. best to start with turing's answer

OpenStudy (turingtest):

yeah I was kinda gonna say...

OpenStudy (anonymous):

just showing off really but if it is to turn in maybe that is what they want to see. or if it is online. who knows.

OpenStudy (anonymous):

Got it, thanks